OpenStudy (ajanijones):

Please help.. medall.. Find the experimental probability. Sheila drew single marbles from a bag, recorded the outcome, and replaced the marble. In 50 draws, Sheila drew 20 white marbles and 30 black marbles. What is the experimental probability that the next marble will be black? A. = 2/5 B. = 3/5 C. = 2/3 D. = 3/7

OpenStudy (ajanijones):

@iceicebaby @mathpapa @mathstudent55 please help

OpenStudy (ajanijones):

Please help if you can..

OpenStudy (anonymous):

to find the probability you want to divide the # black marbles by # total marbles

OpenStudy (ajanijones):

Ok 1 sec.

OpenStudy (ajanijones):

1.5

OpenStudy (anonymous):

sorry my first answer was a typo

OpenStudy (anonymous):

the denominator should be the (# total marbles = 20+30)

OpenStudy (ajanijones):

ok so 50

OpenStudy (anonymous):

yes

OpenStudy (ajanijones):

Then what..

OpenStudy (anonymous):

the formula is (#black)/(#total)

OpenStudy (anonymous):

30/50

OpenStudy (anonymous):

then simplify

OpenStudy (ajanijones):

Ok so you have to simplify

OpenStudy (anonymous):

yes

OpenStudy (ajanijones):

I got B as the answer.?

OpenStudy (anonymous):

looks good to me

OpenStudy (ajanijones):

Thank you i will call you if i need help cause i have 4 more you explain that so well thank you.

OpenStudy (anonymous):

cool, good luck w it

OpenStudy (ajanijones):

Thank you.

OpenStudy (anonymous):

you're welcome :)

OpenStudy (ajanijones):

Are you still here cause i need help.with another one

OpenStudy (anonymous):

ok let's do one more

OpenStudy (ajanijones):

Ok 1 sec.

OpenStudy (ajanijones):

The table shows the results of a spinner with four colored sections. Use the table to answer the question. The spinner is spun 50 more times. Which is the best prediction for the number of times the spinner will stop on white? A. 5 B. 10 C. 25 D. 40 http://static.k12.com/calms_media/media/1416500_1417000/1416594/1/4a6a128174ef4cce8333d1b7887f0e5c9ae95554/FGA_130913_160910.jpg

OpenStudy (anonymous):

cool. first, can you find the probability that it is white

OpenStudy (anonymous):

remember the formula: (#white)/(#total)

OpenStudy (ajanijones):

Ok.

OpenStudy (ajanijones):

So add them all up then divide white by it.

OpenStudy (anonymous):

ya, what does that calculate to

OpenStudy (ajanijones):

0.49

OpenStudy (ajanijones):

So its D.

OpenStudy (ajanijones):

Cause you have to estimate/guess close to it.

OpenStudy (anonymous):

once you have the probability, then you need to multiply by 50

OpenStudy (anonymous):

because you're gonna spin it 50 times

OpenStudy (anonymous):

.49*50

OpenStudy (anonymous):

(probability)*(total spins) = (# of expected spins)

OpenStudy (anonymous):

.49*50 = 24.5

OpenStudy (anonymous):

so pick the one closest to 24.5

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

C. 25 looks like the closest